1. A racing car starts from rest and reaches a final speed v in a time interval of t after traveling a distance on a straight line. If the acceleration a of the car is constant during this time, which of the following statements can not be true?

(a) The car travels a distance of V^2/2a .

(b) The acceleration of the car is v/2t^2 .

(c) The velocity of the car at t/2 (at one half of the time interval t) is 1/2 at^2 .

which of the following statements is true?

A) Statement (a) only

B) Statement (b) only

C) Statement (c) only

D) Statement (b) and (c)
Answer #1

From V^2 - U^2 = 2aS

=) S = V^2/2a

Option a is correct .

Now from V - U = at

=) a = V/t

Velocity at t1/2 ;

From ; V = at/2

Hence statement a is correct only.